OpenStudy (anonymous):

Bill and Amy want to ride their bikes from their neighborhood to school which is 14.4 kilometers away. It takes Amy 40 minutes to arrive at school. Bill arrives 20 minutes after Amy. How much faster (in meters/second) is Amy's average speed for the entire trip?

OpenStudy (anonymous):

Velocity = distance / time Amys velocity = 14.4 / 40 = 0.36 kilometers per minute Bills velocity = 14.4 / (40+20) = 0.24 kilometers per minute 0.36 - 0.24 = 0.12 kilometers per minute

OpenStudy (anonymous):

Thank you! That answered it!